Factorize: a^2 – (b – 5)a -5b : Note it is A square 2

Factorize:
a^2 – (b – 5)a -5b : Note it is A square 2

Grade:

1 Answers

Ravi
9 years ago
a2 – (b – 5)a -5b
OPen the brackets by multiplying the terms in multiplication
=a2-ab+5ab-5b
Take common ‘a’ from first 2 terrms and ‘5b’ from the last 2 terms so as to get the residual ‘(a-b)’same
=a(a-b)+5b(a-b)
Treat (a-b) as an entity and take (a-b) common from the whole expression
=(a-b)[a+5b]
.......................................................................................................................................